OpenStudy (anonymous):

The linear function f(x) has a slope of -2 and intersects the y-axis at the point (0, 14). If g(x) = -2x - 4, which statement is true? The functions f(x) and g(x) do not intersect the x-axis. The function f(x) intersects the x-axis closer to zero than the function g(x). The functions f(x) and g(x) intersect the x-axis at the same point. The function g(x) intersects the x-axis closer to zero than the function f(x).

OpenStudy (gorv):

f(x) is linear implies it is a line as slope is given so f(x)=y=mx+c m and point are given put this in equation

OpenStudy (gorv):

f(x)=y=-2x+c for finding c put (0,14) in it now f(x)=14=-2*0+c implies c=14 f(x)=y=-2x+14 g(x)=-2x-4

OpenStudy (gorv):

at x axis y=0 so to find out at which point it will cut x axis ....put y=0 f(x)=0=-2x+14 -2x=-14 x=7 it will cut x at x=7 g(x)=0=-2x-4 -2x=4 x=-2

OpenStudy (anonymous):

So would it be A

OpenStudy (gorv):

so gx(x) cut x axis closer to zero than f

OpenStudy (gorv):

they are intersecting the x axis

OpenStudy (gorv):

see the steps

OpenStudy (gorv):

D

OpenStudy (anonymous):

yes i understand now thanks